Newdex has net income of $3,000,000 (INCLUDING the effect of expected out-of-pocket costs) and 1,000,000 shares outstanding. It needs to raise $5,000,000 in funds for a new asset. Its investment banker plans to sell an issue of common stock to the public for $40, less a spread of 10%. How much must Newdex's after-tax income be to prevent dilution of earnings per share? Multiple Choice $3,040,000 $3,416,667 $3,350,000 $3,375,000
Answer #1

Current Earnings per share = Net Income/Number of shares

= 3,000,000/1,000,000

= $3 per share

Number of shares required to be issued = 5,000,000/(40 - 10%)

= 138,889 shares

Increase in Income required = New Shares*Current Earnings per share

= 138,889*3

= $416,667

Hence, required after tax income = $3,416,667